Grade 12General Physics

A projectile is fired at an angle of 30o with the horizontal such that the vertical component of its initial velocity is 80 m/s. Find the approximate velocity of the projectile at time T/4 where T is the time of flight.

Mansi
10 Years agoGrade 12

1 Answer

Vikas TU
10 Years ago
Dear Mansi,
.
Using given conditions first:
 usin30’ = 80
u = 80*2 = 160 m/s
Now Calculating the time of flight first,
T  = 2usinthetha/g = 2*160sin30/g
   = 160/g
  = 16 seconds. 
 
at T → T/4
16/4 = 2Vsin30/g
V/g = 4
V = 4g
  = 40 m/s (approx.)using g = 10 m/s^2